OpenStudy (anonymous):

i need help plz Find the side of a square whose diagonal is of the given measure.

OpenStudy (anonymous):



OpenStudy (anonymous):

@I_love_my_nieces

OpenStudy (i_love_my_nieces):

Is that \[15\sqrt{2 cm}\]

OpenStudy (anonymous):

yes

OpenStudy (i_love_my_nieces):

Okay, now do you have a calculator?

OpenStudy (anonymous):

yes i got one

OpenStudy (i_love_my_nieces):

All you do it out 15 then hit the square root button and then type in 2 after the square root and then hit enter and you should get \[15\sqrt{2} = 21.21\]

OpenStudy (anonymous):

yes i got that but is that my answer ?

OpenStudy (i_love_my_nieces):

What are your choices?

OpenStudy (anonymous):

there isn't i have to type it in

OpenStudy (anonymous):

no i do odysseyware

OpenStudy (i_love_my_nieces):

Okay. Do you have to show your work?

OpenStudy (anonymous):

no

OpenStudy (i_love_my_nieces):

Okay, well all you have to put is \[15\sqrt{2 cm} = 21.21 cm\]

OpenStudy (anonymous):

ok thank you and do you have time for more /.\

OpenStudy (i_love_my_nieces):

But MAKE SURE you put cm after the 2 and the 21.21

OpenStudy (i_love_my_nieces):

I'll try :)

OpenStudy (i_love_my_nieces):

Can I have a medal please. And post it as a new question:)

OpenStudy (anonymous):

ok lol and how do i medal doe?

OpenStudy (i_love_my_nieces):

You hit 'Best Response'

OpenStudy (anonymous):

so like i just did lol cx

OpenStudy (i_love_my_nieces):

Yep. Do you still need help?

OpenStudy (anonymous):

yes plz

OpenStudy (i_love_my_nieces):

Ok. Post then in a new post please.

OpenStudy (i_love_my_nieces):

them*

OpenStudy (anonymous):

so i close this question?

OpenStudy (i_love_my_nieces):

Yes.

OpenStudy (anonymous):

ok